The coat color gene of a population of squirrels is defined by a dominant allele G, which encodes a gray coat, and a recessive allele g, which encodes a black coat. In a population, genetic typing reveals the genotypic distribution to be 50 GG, 30 Gg, and 20 gg. If this population is in equilibrium, then what should the genotypic frequency of Gg squirrels in this population be?

0.52

0.46

0.42

0.55
Answer #1

The allele frequency gg q = 0.35

So P+Q = 1 ; P = 1-0.35 = 0.65

The genotypic frequency of Gg = 2PQ = 2 X 0.65 X 0.35 = 0.46